OpenStudy (anonymous):

if polynomial of x=x²-9x-a has 14 and b as its roots then find the value of a & b

OpenStudy (aum):

\(x^2-9x-a = (x-14)(x-b) = x^2 - 14x - bx + 14b = x^2-(14+b)x + 14b\) Therefore, 14 + b = 9 and 14b = -a b = 9 - 14 = -5 and 14(-5) = -a. a = 70 a = 70 b = -5